# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 1.2. Main theorems
We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplicative operators, thereby complementing the result of Le Masson and Sahlsten [LS17]. We also prove a probabilistic analogue of this, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24]. The small difference is that we use the quantum Hamiltonian $\sqrt{-\Delta-1/4}$ instead of $\sqrt{-\Delta}$ but causes only a spectral shift. We already refer to Section 3 for more technical overview of the proof ideas as well as the novel ideas of the proofs used.
The proof follows the ideas of Le Masson and Sahlsten in that we establish a quantitative version which yields limiting forms of the theorem. We again divide the argument into a spectral and a geometric component. However, we employ a different propagator, namely the hyperbolic wave propagator
The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=2^{-1/2}\,P_{t}\,u_{0}$ with initial data
$u(z,0)=0,\qquad u_{t}(z,0)=2\pi f,$
Unlike the ball-averaging operator used by Le Masson and Sahlsten, which has no immediate connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner. This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten.
To access the off-diagonal elements $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we also introduce a modified propagator: $P_{t,\tau}=\cos(t\tau)\,P_{t}$. Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behaviour of the off-diagonal elements. This, however, complicates both the spectral and geometric analyses compared to the quantum ergodicity case, which concerns only the diagonal elements. In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, appearing in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$. The physical interpretation of this operator is not immediately transparent, but it can be verified that
$P_{t,\tau}^{*}aP_{t}=P_{t,\tau/2}^{*}aP_{t,\tau/2}+P_{-t,\tau/2}^{*}aP_{-t,\tau/2},$
meaning it equals the sum of two Heisenberg evolutions of $a$ with $P_{t,\tau/2}$, one forward and one backward in time.
Finally, we replace rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics. The idea of using the exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [HLM+]. The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result. Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum weak mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale. The more technical approach taken in the proofs is presented in Section 3.
We now state the main theorems.
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace-Beltrami operators $(-\Delta_{X_{n}})$. Laplacian $\Delta_{X}$ has eigenvalues
@@
together with a corresponding orthonormal basis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\frac{1}{4}}.$
@@
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplicative operator. Then the following properties hold:
@@
    $\lim_{n\to\infty}\ \frac{1}{\#\{j\ :\ \rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a\,dx\right|^{2}=0,$
@@
Properties (i)(iii) are the large-scale analogues of Properties (i)(iii) in Theorem 1.1. Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying (i)(iii) as large-scale quantum mixing. As in Theorem 1.1, the choice of observables plays a role, recall the example on the unit circle where quantum mixing fails discussed in the earlier remark, Remark 1.3. We note that this idea can be modified to prove that the $2$-torii of increasing area are not quantum mixing for physical-space observables, though there are sequences of $2$-torii that are, interestingly, quantum ergodic using the explicit formula of the eigenfunctions. We refer to the earlier illustrating example discussed This observation also illustrates that large-scale quantum ergodicity and large-scale quantum weak mixing are properties specific to certain sequences of hyperbolic surfaces rather than to general sequences of Riemannian manifolds.
@@
Le Masson and Sahlsten [LS17] have already proven that property (i) holds in this setting. Hence to establish Theorem 1.4, it remains to prove properties (ii) and (iii). To prove the previous theorem, we need the following quantitative version of it.
@@
Let $X$ be a compact connected hyperbolic surface with Laplace-Beltrami operator $\Delta_{X}$ that has eigenvalues
@@
together with a corresponding orthonormal basis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values to eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplicative operator. Let
@@
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I:\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},b\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,8\pi^{2}\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius being smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, such that
@@
We observe from the statement of Theorem 1.5 that in the proof of Theorem 1.4 in Section 8, we could relax many of the conditions. E.g. we could let the spectral gap and the injectivity radius to shrink slowly as we proceed in the sequence. Thereby Theorem 1.4 could be extended to encompass a larger family of sequences.
An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces. This follows from Katz et al. [KSV07], who showed that the injectivity radius grows (ensuring properties (BSC) and (UND)), and from Kim et al. [KSR03], who established a uniform lower bound on the spectral gap (ensuring property (EXP)). The result of Kim et al. is the best deterministic bound currently available. The first, weaker, bound is Selberg’s $3/16$ bound, see [Sel65].
Moreover, the class of admissible sequences is quite broad which is seen when using probabilistic approach using the Weil-Petersson random model; for background, see [Mir13, Wri20]. By Le Masson and Sahlsten [LS24], for a sequence $(X_{g})$ of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is the Weil-Petersson probability measure on the moduli space $\mathcal{M}_{g}$ of hyperbolic surfaces of genus $g$, the condition (i) of Theorem 1.4 holds with high probability as $g$ grows to infinity. We state now a probabilistic analogue of Theorem 1.4, thus showing that a growing Weil-Petersson random surface sequence is quantum ergodic and quantum mixing with high probability and further complementing the results of Le Masson and Sahlsten [LS24].
Theorem 1.7 (Probabilistic version).
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil-Petersson probability distribution on moduli space $\mathcal{M}_{g}$, where $g$ is the genus. Let $\Delta_{g}$ be the Laplace-Beltrami operator of $X_{g}$ with eigenvalues
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values to eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\frac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{g})$ be a sequence of uniformly bounded functions on $(X_{g})$, each acting as a multiplicative operator. Then Properties (i), (ii), and (iii) of Theorem 1.4 hold with high probability as genus $g$ grows to infinity.
As mentioned, Property (i) was proved by Le Masson and Sahlsten [LS24]. Hence it remains to prove Properties (ii) and (iii) where the appraoch of the proof follows closely that of Theorem 1.4. However, as the assumptions are lighter in the probabilistic setting we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law, Theorem 8.1. We need analogues of these in the proof of Theorem 1.7. Theorem 8.2 consisting of recent results in the random surface theory offers us the required tools.
The use of Weil-Petersson random surfaces in quantum ergodicity is currently very active. Randomization effectively suppresses pathological cases helping to establish stong results that the analogous deterministic settings are far away from acquiring. For results in Benjamini-Schramm convergence and injectivity radius in Weyl-Peterson random model, see [Mir13, Mon21]. For recent results on the spectral gap, see [AM24, LW24, Mag24, HMT25, HMN25, WX25]. There are other random models and we especially refer to late groundbreaking results using the random covering model, see [MN20, MNP22, HM23]. See an interesting application of Property (EXP) to the Bose-Einstein condensation of random hyperbolic surfaces [LS22].
@@ 1.3. Further discussions
Like Le Masson and Sahlsten [LS17], we restrict the energy window to a compact interval in $\left(1/4,\infty\right)$. This assumption plays a crucial role in our proofs. However, it is plausible that with a different treatment of the low-lying modes—namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in [0,1/4)[0,1/4)—one could extend the results to an energy window covering the full range $0,\infty)$.
In contrast to Theorem 1.1, Theorem 1.4 is not a characterization-type result linking the classical and quantum dynamics. A natural question arises: can Theorem 1.4 be complemented to yield such a characterization? As noted in Remark 1.6, the sequences of surfaces may in fact possess slowly decaying injectivity radii and spectral gaps. Given the connection between weak mixing and the spectral gap, it is plausible to conjecture that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum weak mixing if and only if the spectral gap tends to zero sufficiently slowly. Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable.
While full characterization results may still be out of reach, the present results appear extendable—albeit with varying levels of difficulty. Several directions for possible extensions present themselves. For instance, can analogous results be established in other structures such as graphs? We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal elements by adapting the techniques developed here, thereby enabling a proof of large-scale weak mixing in the graph setting of Anantharaman and Le Masson. Doing quantum mixing on the graphs seems highly timely as Bordenave, Letroit, and Sabri have an upcoming result on mixing on growing Schreier graphs [BLS].
@@
Another avenue of extensions concerns the choice of observables: in Theorem 1.1, the observable class was quite broad—zeroth-order pseudodifferential operators—so it seems likely that with minor modifications, a larger class of observables could be accommodated in the large-scale setting as well. This has already been achieved in the large-scale quantum ergodicity contexts for graphs [AL15] and surfaces [ABL23], where the kernels of the observables are concentrated near the diagonal but not restricted to it, unlike in the present work.
@@
Yet another natural question is whether the theorem can be adapted to the case of variable curvature, or to accommodate more general energy operators than the Laplacian. For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with general Schrödinger operators, under suitable restrictions. Non-regular graphs can be viewed as discrete analogues of variable-curvature surfaces. It would also be of great interest to construct counterexamples—sequences of surfaces demonstrating that the conditions imposed on the sequences are indeed necessary. In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by MacKenzie [McK22]; however, constructing such examples in the surface setting remains substantially more difficult.
@@
We pose perhaps an obvious question: is it possible to remove the averaging present in our theorems? In the large-energy regime, this corresponds to the quantum unique ergodicity conjecture of Rudnick and Sarnak, see [RS94]. The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces [Lin06]), but remains open in general. Analogous large-scale results are known for graphs; for instance, large-scale quantum unique ergodicity for random regular graphs was proved by Bauerschmidt et al. [BHY19]. We conjecture that the conclusions of Theorem 1.4 remain valid even without averaging. We refer to such statements as large-scale quantum unique ergodic and large-scale quantum unique mixing. These would parallel the averaging-free results of Cipolloni et al. [CES21]. A possibility is to try to establish this first using random models, such as Weil-Petersson model, or by using arithmetic surfaces and thus complement Lindenstrauss [Lin06]. In particular, for arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected via the Watson $L$-function [Wat02].
@@
In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos—such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [Deu18]. Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance of the eigenvalues of $\psi_{j}$ and $\psi_{k}$ goes to infinity. Looking at the large-scale results discussed here, Theorems 1.4 and 1.7, we see that parameter $\tau$ controls the distance of the eigenvalues. However, the eigenvalues are taken from the fixed energy windows which means that for large enough $\tau$, the averages of the off-diagonal elements will be zero as the large distance is not possible in the energy window. However, if we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions. It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will close to zero, but verifying this requires careful study of the sequence of growinf surfaces, $\tau$, and the energy window.

### v2

@@ 1.2. Main theorems
We prove a large-scale analogue of the corollary of Theorem 1.1 for physical-space observables, that is, for functions viewed as multiplication operators, thereby complementing the result of Le Masson and Sahlsten [LS17]. We also prove a probabilistic analogue of this using the Weil–Petersson random model, complementing the probabilistic quantum ergodicity result of Le Masson and Sahlsten [LS24]. A small difference compared to the large-energy setting is that we use the quantum Hamiltonian $\sqrt{-\Delta-\tfrac{1}{4}}$ instead of $\sqrt{-\Delta}$; however, this causes only a spectral shift. We proceed to introduce the main theorems now.
@@
Let $(X_{n})$ be a sequence of compact connected hyperbolic surfaces with corresponding Laplace–Beltrami operators $(-\Delta_{X_{n}})$. The Laplacian $\Delta_{X_{n}}$ has eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(n)}\}$ over $L^{2}(X_{n})$. Define associated values of the eigenvalues as follows:
$\rho_{j}^{(n)}=\sqrt{\lambda_{j}^{(n)}-\tfrac{1}{4}}.$
@@
Let $I\subset(0,\infty)$ be a compact interval and let $(a_{n})$ be a sequence of uniformly bounded functions on $(X_{n})$, each acting as a multiplication operator. Then the following properties hold:
@@
    $\lim_{n\to\infty}\ \frac{1}{\#\{j:\rho_{j}^{(n)}\in I\}}\sum_{j:\rho_{j}^{(n)}\in I}\left|\left\langle\psi_{j}^{(n)},a_{n}\psi_{j}^{(n)}\right\rangle-\frac{1}{|X_{n}|}\int_{X_{n}}a_{n}\,dx\right|^{2}=0,$
@@
Properties (i)(iii) are the large-scale analogues of Properties (I)(III) in Theorem 1.1. Accordingly, we refer to systems satisfying Properties (i) and (ii) as large-scale quantum ergodic, and those satisfying Properties (i)(iii) as large-scale quantum mixing. We remark that Property (i) was proven by Le Masson and Sahlsten in [LS17]. Property (i) is also the surface analogue of the large-scale quantum ergodicity theorem of Anantharaman and Le Masson [AL15] discussed in the previous subsection.
In Appendix A, we construct an example of a sequence of growing flat $2$-tori for which Properties (ii) and (iii) fail. This demonstrates that large-scale quantum mixing is not a generic property for sequences of Riemannian surfaces. It remains an interesting question whether Property (i) holds for sequences of growing tori; establishing this would complement the large-energy result, analogous to Property (I) of Theorem 1.1, proved by Marklof and Rudnick [MR12]. For the standard eigenfunctions
$\psi_{(m,n)}(x,y)=e^{-2\pi i\left(\frac{mx}{L_{x}}+\frac{ny}{L_{y}}\right)},\qquad(m,n)\in\mathbb{Z}^{2},$
the claim is straightforward. However, establishing the result for all eigenbases seems to be subtle.
An example of a sequence of surfaces that is both quantum ergodic and quantum mixing in the above sense is a sequence of compact arithmetic surfaces. This follows from the result of Katz, Schnaps, and Vishne [KSV07], who showed that the injectivity radius grows (ensuring Properties (BSC) and (UND)), and from the result of Selberg [SEL65], who established a uniform lower bound of $\tfrac{3}{16}$ on the spectral gap (ensuring Property (EXP)). We remark that the result of Kim, Sarnak, and Ramakrishnan [KRS03] is the best deterministic bound currently available on the spectral gap.
Moreover, the class of admissible sequences is quite broad, which is seen when using the probabilistic approach via the Weil–Petersson random model; for background on the Weil–Petersson random model, see [MIR13, WRI20]. We state the probabilistic analogue of Theorem 1.3, thus showing that a sequence of growing Weil–Petersson random surfaces is quantum ergodic and quantum mixing with high probability.
Theorem 1.4 (Probabilistic version).
Let $(X_{g})$ be a sequence of $\mathbb{P}_{g}$-random surfaces, where $\mathbb{P}_{g}$ is a Weil–Petersson probability distribution on the moduli space $\mathcal{M}_{g}$, and where $g$ is the genus. Let $\Delta_{g}$ be the Laplace–Beltrami operator of $X_{g}$ with eigenvalues
together with a corresponding orthonormal eigenbasis $\{\psi_{j}^{(g)}\}$ over $L^{2}(X_{g})$. Define associated values of the eigenvalues as follows:
$\rho_{j}^{(g)}=\sqrt{\lambda_{j}^{(g)}-\tfrac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval, and let $(a_{g})$ be a sequence of uniformly bounded functions on $X_{g}$, each acting as a multiplication operator. Then Properties (i)(iii) of Theorem 1.3 hold with high probability as the genus $g$ grows to infinity.
We remark that the use of randomization is highly active in the field of quantum chaos. Randomization effectively suppresses pathological cases, helping to establish strong results; the analogous deterministic settings are far from acquiring the corresponding results. We especially want to highlight the following important results: the Benjamini–Schramm convergence and a lower bound on the injectivity radius in the Weil–Petersson random model were established in [MIR13, MON21]. For recent results on the spectral gap, see [AM24, LW24, MAG24, HMT25a, HMN25b, WX25]. Another random model, the random covering model, has been used with great success to obtain results on the spectral gap; see [MN20, MNP22, HM23].
@@
As mentioned, Le Masson and Sahlsten have already proven that Property (i) holds in the setting of Theorem 1.3 (see [LS17]). They also proved that Property (i) holds in the setting of Theorem 1.4 (see [LS24]). Hence, to establish Theorems 1.3 and 1.4, it remains to prove Properties (ii) and (iii) in both settings.
The approaches used to prove Theorems 1.3 and 1.4 are very similar: they rely on the quantitative version of the theorems, namely Theorem 1.5. Using the quantitative theorem together with Properties (BSC), (EXP), (UND), and the large-scale Weyl law (see Theorem 8.1) yields a relatively straightforward proof of Theorem 1.3.
When proving the probabilistic theorem, Theorem 1.4, we do not have Properties (BSC), (EXP), and (UND) available, nor do we have the large-scale Weyl law. These tools can be replaced with those recorded in Theorem 8.2. We remark that Theorem 8.2 is compiled from [LS24, Thm. 1.7, 1.9]; the references given for this theorem differ from those of Le Masson and Sahlsten [LS24], reflecting that the understanding of the spectral gap is evolving rapidly.
Next, we state the quantitative main theorem and comment on the approach taken in its proof.
@@
Let $X$ be a compact connected hyperbolic surface with Laplace–Beltrami operator $\Delta_{X}$ that has eigenvalues
@@
together with a corresponding orthonormal eigenbasis $\{\psi_{j}\}$ over $L^{2}(X)$. Define associated values of the eigenvalues as follows:
$\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}.$
Let $I\subset(0,\infty)$ be a compact interval and let $a$ be a mean 0 function on $X$ acting as a multiplication operator. Let
@@
$\sum_{j:\rho_{j}\in I}\ \sum_{\begin{subarray}{c}k:\rho_{k}\in I,\\ k\neq j,\\ |\rho_{j}-\rho_{k}-\tau|<\delta\end{subarray}}\left|\langle\psi_{j},a\,\psi_{k}\rangle\right|^{2}\leq C_{8}\,\max(I)^{4}\,\frac{1}{T\,\beta^{3}}\,\left(\|a\|_{2}^{2}\,+\,\|a\|_{\infty}^{2}\,|X\setminus X(4T)|\frac{e^{6T}}{\operatorname{InjRad}_{X}}\right),$
where $|X\setminus X(4T)|$ is the volume of the points of $X$ with injectivity radius smaller than $4T$, and where $\beta=\beta(\lambda_{1})$, defined by
@@
We observe from the statement of Theorem 1.5 that we could relax many of the conditions appearing in the statement of Theorem 1.3. For example, we could let the spectral gap and the injectivity radius slowly shrink as we proceed along the sequence. Thereby, Theorem 1.3 could be extended to encompass a larger family of sequences of surfaces.
The proof of the quantitative theorem also follows the idea of Le Masson and Sahlsten [LS17] in that we divide the argument into a spectral and a geometric component. However, we employ a different propagator, namely the hyperbolic wave propagator
The intuition behind $P_{t}$ is that $u_{tt}=-\Delta\,u$, the wave equation in the hyperbolic plane, has solution $u=P_{t}\,f$ with initial data
$u(z,0)=0,\qquad u_{t}(z,0)=f,$
Unlike the ball-averaging operator used by Le Masson and Sahlsten [LS17], which does not have as immediate a connection to the underlying classical dynamics, the hyperbolic wave propagator is obtained from the quantization of the geodesic flow, thus linking the quantum and classical settings in a natural manner. This choice of propagator renders the spectral side particularly clean, though it makes the geometric side more involved compared to the approach of Le Masson and Sahlsten [LS17].
To access the off-diagonal averages $\langle\psi_{j},a\psi_{k}\rangle$, where the energy levels of $\psi_{j}$ and $\psi_{k}$ may differ substantially, we introduce a modified propagator
$P_{t,\tau}=\cos(t\tau)\,P_{t}.$
Using $P_{t}$ and $P_{t,\tau}$ simultaneously allows us to capture the behavior of these off-diagonal averages. However, this complicates both the spectral and geometric analyses compared to the diagonal case studied by Le Masson and Sahlsten [LS17].
In particular, the Heisenberg evolution of the observable $a$, $P_{t}^{*}aP_{t}$, which appears in the proof of quantum ergodicity by Le Masson and Sahlsten, is replaced here by $P_{t,\tau}^{*}aP_{t}$. The physical interpretation of this operator is not immediately transparent, but it can be formally verified that
$P_{t,\tau}^{*}aP_{t}=\cos(t\tau)\,P_{t}^{*}aP_{t}=2\cos\!\left(t\frac{\tau}{2}\right)P_{t}^{*}a\cos\!\left(t\frac{\tau}{2}\right)P_{t}-P_{t}^{*}aP_{t}=2P_{t,\frac{\tau}{2}}^{*}aP_{t,\frac{\tau}{2}}-P_{t}^{*}aP_{t},$
where we used the identity $\cos(2x)=2\cos^{2}(x)-1$. This shows that $P_{t,\tau}^{*}aP_{t}$ can be written as the subtraction of two Heisenberg evolutions of $a$, one with $P_{t,\frac{\tau}{2}}$ and one with $P_{t}$.
Finally, we replace the rather abstract Nevo’s theorem with the exponential mixing theorem (see Theorem 1.2), which elucidates the connection between classical and quantum dynamics. To use the exponential mixing theorem, we open the squares of the Hilbert–Schmidt norm appearing in the proof, as opposed to Le Masson and Sahlsten [LS17], who opt to use Minkowski’s integral inequality. In this way, we lose less decay and see an interesting exponential decorrelation of Heisenberg evolutions $P_{t}^{*}aP_{t}$ and $P_{t^{\prime}}^{*}aP_{t^{\prime}}$ when $|t-t^{\prime}|$ increases. The idea of using exponential mixing originates in the upcoming article of the author with Lequen, Mikkelsen, Sahlsten, and Ueberschär [30]. The use of exponential mixing is both simple and transparent: the exponential decay of correlations of $a$ and $a\circ\varphi_{t}$, even for small $t$, is essential in obtaining our result. Small propagation times suffice because we work with observables in a macroscopic sense, unlike the large-energy quantum ergodicity and quantum mixing setting, which employs microlocal analysis and requires long propagation times for decorrelation to occur at the microlocal scale. We already refer to Section 3 for a more technical overview of the proofs.
@@ 1.3. Further discussions
Like Le Masson and Sahlsten [LS17], we restrict the energy window to a compact interval in $\left(1/4,\infty\right)$. This assumption plays a crucial role in our proofs. However, it is plausible that with a different treatment of the low-lying modes, namely, those eigenvalues and corresponding eigenfunctions with eigenvalues in $[0,1/4]$, one could extend the results to an energy window covering the full range $0,\infty)$.
@@
Another avenue of extension concerns the choice of observables. In Theorem 1.1, the observable class was quite broad, zeroth-order pseudo-differential operators, so it seems that with some modifications, a larger class of observables could be accommodated in the large-scale setting as well. In the large-scale quantum ergodicity setting for graphs [AL15] and surfaces [ABL23], a broader class of observables is already considered when treating the diagonal averages (Property (i)): namely, observables whose integral kernels are concentrated near the diagonal, rather than being strictly supported on it as in the present work.
In contrast to Theorem 1.1, Theorem 1.3 is not a characterization-type result linking the classical and quantum dynamics. A natural question arises: can Theorem 1.3 be complemented to yield such a characterization? As noted in Remark 1.6, the sequences of surfaces may in fact possess slowly decaying injectivity radii and spectral gaps. Given the connection between weak mixing and the spectral gap, it is plausible that a sequence of hyperbolic manifolds converging to a hyperbolic space in the Benjamini–Schramm sense is large-scale quantum mixing if and only if the spectral gap tends to zero sufficiently slowly. Establishing such characterizations for both large-scale quantum ergodicity and large-scale quantum mixing would be remarkable for further illustrating the connections between the classical and quantum settings.
While full characterization results may still be out of reach, several directions for possible extensions present themselves, albeit with varying levels of difficulty. For instance, can analogous results be established in other structures such as graphs? We conjecture that the graph result of Brooks, Lindenstrauss, and Le Masson [BLL16] could be extended to include off-diagonal averages in the setting of the large-scale quantum ergodicity theorem for graphs by Anantharaman and Le Masson [AL15] by using the ideas presented here. Performing quantum mixing on graphs seems highly timely, as Bordenave, Letrouit, and Sabri have a recent result on quantum mixing on growing Schreier graphs [BLS26].
@@
Yet another natural question is whether the results can be adapted to the case of variable curvature or to accommodate more general energy operators than the Laplacian. For graphs, analogous results are known: Anantharaman and Sabri [AS17, AS19] established large-scale quantum ergodicity for non-regular graphs with more general Schrödinger operators than the Laplacian, under suitable restrictions. Non-regular graphs can be viewed as discrete analogues of variable-curvature surfaces. It would also be of great interest to construct counterexamples that would highlight the necessity of the conditions imposed on the sequences, namely Properties (BSC), (EXP), and (UND). In the graph setting, the results of Anantharaman and Le Masson were complemented in this manner by McKenzie [MCK22]; however, constructing such examples in the surface setting seems substantially more difficult.
@@
We pose perhaps an obvious question: is it possible to remove the averaging present in our theorems? In the large-energy regime, this corresponds to the quantum unique ergodicity conjecture of Rudnick and Sarnak, see [RS94]. The conjecture has been resolved in certain special cases (notably by Lindenstrauss for arithmetic surfaces and Hecke eigenfunctions [LIN06]), but remains open in general. We expect that the conclusions of Theorem 1.3 and Theorem 1.4 remain valid even without averaging, possibly requiring only minor adjustments to the statements. Especially, the averaging-free version of the probabilistic Theorem 1.4 would parallel the averaging-free results for random graphs [BHK+17a, BHY19, BKY17b] and random matrices [CES21]. However, the averaging-free version of the deterministic Theorem 1.3 would more closely parallel the quantum unique ergodicity conjecture; it is likely harder to establish than the probabilistic version.
For arithmetic surfaces, averaging-free off-diagonal bounds are of significant interest: strong estimates on $|\langle\psi_{j},\psi_{k}\psi_{l}\rangle|$ could contribute to a proof of the Lindelöf hypothesis, to which quantum chaos is elegantly connected; see Watson’s work on $L$-functions [WAT02].
@@
In any case, a deeper understanding of averaging-free results may provide a pathway toward results in many-body quantum chaos, such as the Eigenstate Thermalization Hypothesis proposed by Deutsch in the physics literature [DEU18]. Deutsch predicts that $|\langle\psi_{j},a\psi_{k}\rangle|$ goes to zero as the distance between the eigenvalues of $\psi_{j}$ and $\psi_{k}$ increases. Looking at the large-scale results discussed here, Theorems 1.3 and 1.4, we see that the parameter $\tau$ controls the distance of the eigenvalues. However, the eigenvalues are taken from fixed energy windows, which means that for large enough $\tau$, the averages of the off-diagonal terms will be zero, as the large distance is not possible in the energy window. If we let the energy window grow, then $\tau$ can also grow without giving trivial conclusions. It is natural to assume that in this setting of growing energy window and growing $\tau$, the average will be close to zero, but verifying this requires careful study of the sequence of growing surfaces, the parameter $\tau$, and the energy window.